OpenStudy (anonymous):

Consider the function f(x) = 5/x^4 - 8/x^8, let F(x) be the anti derivative of f(x) with F(1) = 0. Then F(4) =?

OpenStudy (jdoe0001):

haven't done much on that amistre64 or maybe Hero would though

OpenStudy (loser66):

what does "anti derivative" mean?

OpenStudy (loser66):

It means F(x) ' = f(x) , right?

OpenStudy (anonymous):

yeah except its f'(x) = f(x)

OpenStudy (loser66):

nope, F'(x) = f(x) not f'(x) = f(x) f'(x) =f(x) if and only if f(x) = e^x

OpenStudy (anonymous):

oh yeah your right

OpenStudy (loser66):

Now, find F(x) you have f(x) = \(5x^{-4}-8x^{-8}\) do the first term first. \((something)'= 5x^{-4}\), can you find that "something" Or, do you study integral yet?

OpenStudy (anonymous):

we've studied integrals

OpenStudy (loser66):

oh, it is wwwwwwwwwway easier, just take integral of f(x) , you get F(x)

OpenStudy (anonymous):

would it be -5/3(x)^-3

OpenStudy (loser66):

do both them at the same time since you are allowed using integral. :)

OpenStudy (anonymous):

and then theres 8/7(x)^-7

OpenStudy (anonymous):

right?

OpenStudy (loser66):

yes

OpenStudy (loser66):

and +C

OpenStudy (anonymous):

right so then i find C?

OpenStudy (loser66):

yep

OpenStudy (anonymous):

im not sure if its right, i got -59/21, can you check it?

OpenStudy (loser66):

nope

OpenStudy (anonymous):

you substitute 1 for x and equal it to 0 right?

OpenStudy (loser66):

\(F(x) = \dfrac{-5}{3}x^{-3}+\dfrac{8}{7}x^{-7}+C\)

OpenStudy (loser66):

now F(1) = 0 that is \(C= \dfrac{5}{3}-\dfrac{8}{7}=\dfrac{35-24}{21}=\dfrac{11}{21}\)

OpenStudy (loser66):

I think you mess up at the second term . Integral of \(-8x^{-8} \) is 8/7 x^(-7)

OpenStudy (anonymous):

i did get that

OpenStudy (loser66):

how about C? To me, I isolate C first, then replace x =1 to find C and it is 11/21 as shown

OpenStudy (anonymous):

yeah i get that

OpenStudy (loser66):

Good, then just plug that C into F(x) to find F(4)

OpenStudy (anonymous):

ok thnx, lets see wat i get?

OpenStudy (anonymous):

ok i got the answer, thnx for ur help @Loser66

OpenStudy (loser66):

yw